# 3cm3cm2cm2cm Checking the Be anomaly with a two-arm electron positron pair spectrometer

### 22 Jan 2024 v1

@@ 4 Summary
We successfully built a two-arm ee spectrometer in Hanoi. The spectrometer was tested and calibrated using the 17.6 MeV M1 transition excited in the Li(p,ee)Be reaction. We have got a nice agreement between the experimentally determined acceptance of the spectrometer with the one coming from our simulation. The angular correlation of the ee pairs measured for the 17.6 MeV transition agrees well with the simulated one for the M1 transition. No anomaly was observed for this, or the = 800 keV transition either. However, a significant anomaly () was observed for the 18.15 MeV transition at around , in agreement with the ATOMKI results published in 2016[1]. The mass of the hypothetical particle from this result was obtained to be (statistical) (systematic) MeV. In the future, we are planning to upgrade the spectrometer to get a wider angular acceptance.

### 19 Mar 2024 v2

@@ 4 Summary
We successfully built a two-arm ee spectrometer in Hanoi. The spectrometer was tested and calibrated using the 17.6 MeV M1 transition excited in the Li(p,ee)Be reaction. We have got a nice agreement between the experimentally determined acceptance of the spectrometer with the one coming from our simulation. The angular correlation of the ee pairs measured for the 17.6 MeV transition (E=441keV) agrees well with the simulated one dominated by the M1 transition, and no anomaly was observed for this. However, a significant anomaly () was observed for E=1225 keV, which is the off resonance region above the E=1040 keV resonance, at around 135, in agreement with the ATOMKI results published in 2016[1]. The mass of the hypothetical particle from this result was obtained to be (statistical) (systematic) MeV. In the future, we are planning to upgrade the spectrometer to get a wider angular acceptance.